OpenStudy (anonymous):

just need to check ans. plse: Add simplify by collecting like radical terms if possible assuming all expressions under radicals represent a nonnegative numbers: √5a + 3√125a^3 = (1 + 15a)√5a is my ans. plse check

OpenStudy (anonymous):

is it cubed root or square root?

OpenStudy (anonymous):

\[\sqrt{5a}+3\sqrt{125a^3}=\sqrt{5a}+5a\sqrt{5a}=(1+5a)\sqrt{5a}\]

OpenStudy (anonymous):

not sure if your 15 was a typo or not